### v2

# Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling

Abstract
Creating engaging narratives from visual data is crucial for automated digital media consumption, assistive technologies, and interactive entertainment. This survey covers methodologies used in the generation of these narratives, focusing on their principles, strengths, and limitations. The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering. These tasks share common challenges with Visual Story Generation (VSG) and have served as inspiration for the techniques used in the field. We analyze the main datasets and evaluation metrics, providing a critical perspective on their limitations.
keywords:
Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling
[…]
1 Introduction
In this document we survey the field of story generation from visual inputs, covering techniques, related tasks, and challenges. Stories are fundamental to human experience, serving as a bridge between imagination and reality Hühn et al. (2014). They represent the art of telling tales, whether real or imagined, and present a sequence of events that can captivate, inform, and provoke thought. Storytelling has been an essential part of human culture for millennia, serving as a means of communication, entertainment, education, and moral guidance. The art of storytelling has evolved over time, beginning with the practice of narrating stories through spoken words, advancing to written literature, and, more recently, transitioning to digital media formats. Stories have played a crucial role in shaping societies, passing down knowledge and values from one generation to another. Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology Abbott (2008). The field of story generation has gained attention, fueled by advances in Natural Language Processing (NLP) Jurafsky and Martin (2009) and computer vision Szeliski (2010). The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans Gervas (2009). With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data Krishna et al. (2017). We refer to this task as VSG.
1.1 Definition and challenges
VSG is the task of automatically creating coherent and engaging stories grounded on visual inputs, such as images or videos. It extends beyond simple image captioning or video description, requiring an understanding of complex relationships, temporal progression, and implicit context within the visual data. The generated narrative must, not only accurately represent the visual content, but also encapsulate the essence of the underlying story, thereby crafting an engaging experience for the audience. VSG serves as a robust indicator of machine comprehension of visual content and poses several challenges, some of which are:
  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives Xu et al. (2015).
  • Temporal coherence: A story is inherently sequential and maintains a temporal order. Therefore, generated narratives should respect time progression, even if not in a linear manner (e.g., they may involve flashbacks, foreshadowing, or other such mechanisms), and ensure that events are narrated in a logically consistent manner.
  • Contextual understanding: Often, elements of a story can be implicit or inferred from the context. Effective generation approaches must be capable of understanding and incorporating implicit information in the narrative.
  • Engagement and creativity: One of the main goals of a story is to engage the audience. Therefore, the generated narrative should not only be accurate and coherent, but also creative and engaging.
  • Evaluation: Determining the effectiveness of a VSG system is a challenge in itself. Metrics originally developed for tasks such as machine translation and image captioning have been adapted for VSG. However, these metrics are not appropriate to capture the creativity and engagement aspects of a story.
1.2 Background and Motivation
The motivations for VSG are numerous. For example, VSG can provide an interactive component to media consumption. Users could potentially influence the direction of the story by choosing different sequences of images or videos, creating a unique, personalized narrative. This could be particularly beneficial in areas like interactive gaming and movies, educational platforms, or even in digital marketing, where personalized narratives can greatly enhance user engagement and satisfaction. VSG can also be used in journalism: it can be used to summarize large amounts of visual data, giving insights into the underlying patterns and relationships, and assisting in content creation.
The growing interest in this area is also motivated by its potential to contribute to the understanding of human cognition and creativity. By developing models that emulate human-like storytelling abilities, researchers can gain insights into the cognitive processes involved in story creation, as well as explore the underlying mechanisms of creativity. These approaches can further lead to novel applications, such as more engaging virtual experiences Fan et al. (2018); Peng et al. (2018).
These use cases represent only a fraction of the potential applications of VSG. As research in this field progresses, we anticipate that more applications will emerge, spanning an even broader range of domains.
1.3 Scope and structure of the survey
We start by defining the core elements of a story in Section 2 as they are essential for understanding the structure and dynamics of a story. Section 2 presents these elements and discusses their significance in storytelling. It also clarifies the distinctions between story, narrative, and plot, illustrating how a single story can be narrated in multiple ways to yield diverse experiences.
In Section 3 we then explore the main datasets used in story generation, and the various metrics used to assess their performance. We address the limitations that current evaluation metrics face in such an open-ended task and discuss potential directions for future research in this regard.
The connection of VSG with computer vision tasks such as Visual Question Answering (VQA) and image/video captioning is also acknowledged, since they share core challenges, such as visual understanding and narrative generation. Thus, they have contributed with useful techniques to the field. Section 4 explores this convergence between computer vision and NLP.
Concerning VSG proper, we aim to present a balanced perspective on current methodologies, emphasizing opportunities and challenges for future research. The inclusion of a detailed analysis of deep learning methodologies with a broader review of additional and earlier techniques ensures broad coverage of the field. Section 5 presents a broad perspective on the state-of-the-art of VSG.
Section 6 examines the real-world applications of VSG. Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
2 The Elements of Stories
Stories connect events and experiences through narration, ranging from personal tales to cultural epics. While the terms story, narrative, and plot are often used interchangeably, they have distinct meanings: a story is the chronological sequence of events (the “what”), the narrative is how these events are presented (the “how”), and the plot is the specific sequence as experienced by the audience Simpson (2002).
A single story can be told through multiple narratives, each offering different perspectives and experiences. Narratives may follow linear structures with chronological progression, or nonlinear approaches using flashbacks, flash-forwards, or parallel storylines that create complexity and require audience interpretation Lau and Chen (2008); Mitchell (2008).
Story elements – character, conflict, theme, setting, plot, and mode – form the foundation for crafting and analyzing narratives. Understanding these elements is crucial for both human storytellers and automated story generation systems. In the following subsections, we detail each element and its significance in VSG.
2.1 Character
Characters are the entities through which stories unfold, with their decisions and actions propelling the plot forward DiBattista (2011). Through actions and speech, characters convey motives and emotions that shape audience perception and emotional connection. Initial character impressions influence audience expectations about future behavior. Characters who deviate from established patterns can create surprise or confusion, challenging audience understanding. Stories center around protagonists who encounter central conflicts, undergo personal growth, or acquire new knowledge. Antagonists add complexity and tension, often opposing the protagonist’s goals. Traditional narratives feature heroes battling for moral causes against villains who perpetrate evil. Character development and evolution throughout the narrative engages audiences and drives story progression. The range of character types contributes to story depth and dynamics.
2.2 Conflict
Conflict generates tension and drives character actions, shaping narrative trajectory and audience engagement Abbott (2008). How conflicts are introduced, developed, and resolved determines story progression and emotional depth. Conflict resolution provides closure to the central problem and marks the culmination of narrative tension. Conflict encompasses any tension influencing characters, from the protagonist’s main challenge to secondary internal struggles like anxiety or indecisiveness. These layers add complexity to character journeys and story depth. Longer narratives involve multiple conflicts that occur alongside the primary tension. Common conflict types include: Character vs. Character (opposition), Character vs. Nature (natural forces), Character vs. Society (social norms), Character vs. Unavoidable Circumstances (fate), and Character vs. Self (internal dilemmas).
2.3 Theme
Themes are the major underlying ideas of a story, often abstract and open to interpretation Griffith (2010). Unlike concrete story elements, themes invite discussion and allow for multiple interpretations and conclusions. They function as exploratory concepts that encourage audience reflection on deeper meanings.
Themes can evolve as narratives unfold, revealing new dimensions and allowing audiences to develop personal interpretations of the story’s messages. Different audience members may reach varying conclusions that differ from or exceed the creator’s original intent. Understanding themes in narrative generation creates dynamic, thought-provoking storytelling experiences.
2.4 Setting
Setting encompasses the time, place, and context in which the story unfolds Truby (2007). It includes physical and temporal surroundings, as well as social or cultural conventions that influence character actions and decisions. Setting can act dynamically, possessing specific traits and undergoing changes that affect the plot. This allows it to evolve throughout the story, reflecting and impacting narrative progression. Hogwarts School in the “Harry Potter” series exemplifies an evolving setting that influences plot and character development Rowling (2015).
2.5 Plot
Plot is the sequence of events that propels the story from beginning to end Dibell (1999). Each plot point represents a moment of change that influences character understanding, decisions, and actions. This progression typically involves protagonists encountering conflict and attempting resolution, creating emotional stakes and suspense for both characters and audience. Plot advancement deepens audience investment in character fate. Stories can have multiple plots representing different event sequences or perspectives. Plot structure follows cause and effect, where character actions and events lead to reactions and consequences, forming a chain that moves the story forward.
2.6 Mode
Mode encompasses the choices and techniques authors use to frame narratives Pinault (1992). A single story can be narrated using different modes, each offering distinct audience perspectives and experiences. Mode includes information scope, language style, medium, and the extent of narrative exposition and commentary. A key component is point of view or perspective. First-person narratives use personal pronouns like “I” and “me,” directly engaging audiences with the character’s perspective. Third-person narratives avoid personal pronouns, offering more detached viewpoints. These perspectives exemplify narration techniques that explicitly tell stories through spoken or written commentary. While narration is common in written narratives, it can be optional in visual or interactive media.
3 Benchmark Datasets and Evaluation Metrics
This section presents and discusses benchmark datasets used for story generation and evaluation metrics that assess the performance of story generation models. Table 1 provides a compilation of the datasets documented in the literature for VSG. Among these datasets, VIST Huang et al. (2016) is prominently utilized by the research community. This dataset serves as a primary benchmark, and most publications in this field report their experimental results on it. The VIST-Edit dataset Hsu et al. (2019) provides human edited versions of machine generated stories, offering insights into human preferences for story quality. The Visual Writing Prompts dataset Hong et al. (2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images. These datasets allow for the exploration of more complex and imaginative narrative structures, serving as a resource to train models to generate fictional stories.
Dataset No. of Stories Notes
NY City Park and Kim (2015) 11,863 11,863 blog posts / 78,467 images
Disneyland Park and Kim (2015) 7,717 5 7,717 blog posts / 60,545 images
ROCStories Mostafazadeh et al. (2016) 100,000 5 sentences/story.
LSMDC Rohrbach et al. (2016) 118,114 202 movies.
VIST Huang et al. (2016) 50,000 210,819 unique photos
VIST-Edit Hsu et al. (2019) 14,905 Human-edited versions of 2,981 machine-generated stories
Visual Writing Prompts Hong et al. (2023) 12,000 5 to 10 images/story, grounded on the characters.
Pororo-SV Kim et al. (2017) 27,328
YouCook2 Das et al. (2013) 2,000 Average length of 5.26 mins.
CMU movie summary corpus 42,306 Includes box office revenue, genre, release date, runtime, character names and actor names.
Table 1: Visual Story Generation Datasets.
The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures of the quality of generated narratives. Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation or summary, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation. These metricsa include Bilingual Evaluation Understudy (BLEU) Papineni et al. (2002), Self-BLEU Zhu et al. (2018), Metric for Evaluation of Translation with Explicit ORdering (METEOR) Banerjee and Lavie (2005), Consensus-based Image Description Evaluation (CIDEr) Vedantam et al. (2014), Recall-Oriented Understudy for Gisting Evaluation (ROUGE) Lin (2004), Perplexity Jurafsky and Martin (2023) and Bert Score Zhang et al. (2019). These metrics guide the development and refinement of VSG models by offering insights into their linguistic fluency, textual diversity, and mostly lexical overlap with reference texts. They focus on specific linguistic and content-related aspects, providing quantitative measures for precision, recall, and consensus. However, they are limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
Huang et al. Huang et al. (2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics. The results in Table 2 show that METEOR, which places a strong emphasis on paraphrasing, has higher correlation with human judgment than BLEU. Nevertheless, in spite of METEOR showing the highest correlation with human judgment, this value is still low, suggesting that a large gap still exists between automatic metrics and human judgment.
Metric METEOR BLEU
Pearson 0.22 (2.8e-28) 0.08 (1.0e-06)
Spearman 0.20 (3.0e-31) 0.08 (8.9e-06)
Kendall 0.14 (1.0e-33) 0.06 (8.7e-08)
Table 2: Correlations of automatic scores against human judgements, with p-values in parentheses Huang et al. (2016).
To evaluate VSG models, additional metrics that capture the creative aspects of storytelling may be necessary. These metrics might assess the uniqueness, novelty, and diversity of generated narratives, ensuring that models produce stories that go beyond mere data memorization. Coherence metrics could delve deeper into the logical flow of events, ensuring that generated narratives maintain a consistent and plausible storyline.
The proliferation of diverse Large Language Models (LLMs), such as the GPT OpenAI (2024, 2024a, 2024b, 2024c), Llama Touvron et al. (2023); Meta (2024), Mistral team (2023); Mistral AI (2024), and Claude Anthropic (2025, 2024a, 2024b, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks. These scoring systems include Multi-turn Benchmark (MT-Bench) Zheng et al. (2023) and Chatbot Arena Zheng et al. (2023) briefly described below. While they were initially developed for the evaluation of chatbots by human evaluators, they have been adapted to operate using a LLM as the evaluator achieving results close to those obtained by human evaluators.
As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al. (2023). Metrics to evaluate stories could be developed based on LLMs Zheng et al. (2023), possibly following the underlying principles discussed above.
MT-Bench is a benchmark tailored to test the multi-turn conversation and instruction-following capabilities of LLM-based chat assistants. It includes 80 high-quality multi-turn questions distributed across diverse categories such as writing, roleplay, reasoning, and knowledge domains. In the MT-Bench evaluation process, the user is presented with two distinct conversations, each generated by a different LLM-based chat assistant and is then tasked with deciding which assistant, A, B, or indicating a tie, better followed the instructions and answered the questions. For VSG evaluation, this framework could be adapted by using an LLM as an automated judge to compare stories generated by different models, evaluating them based on criteria such as narrative coherence, visual grounding, and creativity.
Chatbot Arena introduces a crowd-sourcing approach, fostering anonymous battles between LLM-based chatbot models. Users interact with two anonymous models simultaneously, posing identical questions to both and subsequently voting for the model that provides the preferred response. The Elo rating system Zheng et al. (2023) is employed to rank the models based on their performance in these battles. This rating system assigns a numerical score to each model based on their battle outcomes, adjusting these scores higher or lower after each encounter to reflect their relative ability to satisfy user queries effectively compared to their opponents. A similar automated evaluation system for VSG could employ an LLM judge to assess pairs of generated stories given the same image sequence, ranking models based on story quality across multiple dimensions without requiring human annotators.
In summary, while current metrics provide insights into specific dimensions of VSG, the development of new metrics may be necessary to holistically evaluate creativity, coherence, and emotional engagement, ensuring a more comprehensive understanding of the capabilities and limitations of these models. We anticipate that LLMs can play a role in the development of such metrics, serving as automated judges that evaluate stories based on multiple criteria.
4 From NLP to computer vision
This section provides an overview of areas at the intersection of computer vision and Natural Language Processing (NLP) that are relevant to the topic of VSG, namely image and video captioning and Visual Question Answering (VQA). These tasks share a common challenge: they require the understanding of visual content and the ability to generate text based on that visual content. Some of the techniques used in these tasks have been adapted to VSG.
4.1 Image and Video Captioning
The transition from image and video captioning to VSG represents a shift from describing individual images or video segments to narrating image sequences or video segments. While the former focus on what is in images or videos, the latter also considers why and how, capturing the underlying narrative that links a series of images or video frames.
VSG models need to understand the temporal progression of events, causal relationships, and the ability to construct cohesive and engaging stories. VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content. Therefore, they build upon the foundational principles of image and video captioning, incorporating both the interpretation of visual data and the generation of coherent, contextually appropriate narratives.
The tasks of image and video captioning consist in generating descriptive, human-readable sentences that represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language.
One notable contribution in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach Vinyals et al. (2014); Yao et al. (2015); Anderson et al. (2017). This technique employs Convolutional Neural Networks (CNNs) as encoders to extract visual features. Specifically, 2D-CNNs are employed for image processing where the model interprets two dimensions: height and width. On the other hand, for video processing, 3D-CNNs are typically used where an additional dimension, time or depth, is taken into account to perceive the temporal dynamics in video sequences. Then, Recurrent Neural Networks (RNNs) or their improved variants, Long Short-Term Memory networks (LSTMs) Hochreiter and Schmidhuber (1997) and Gated Recurrent Units (GRUs) Chung et al. (2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs. This combination of CNNs and RNNs effectively bridges the gap between vision and language.
In the realm of image captioning, Transformers Vaswani et al. (2017) have been used as encoders to process visual features extracted from images. Leveraging the self-attention mechanism inherent in Transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies Olimov et al. (2021); Radford et al. (2021); Liu et al. (2021); Yu et al. (2019). Transformer models also demonstrate the ability to generate precise and elaborate textual descriptions, significantly improving the overall quality of image captions.
Similarly, Transformers have made contributions to video captioning, addressing the challenges associated with understanding and describing temporal dynamics in videos. By treating video frames as a sequential series of images, Transformers can leverage their attention mechanism to capture long-range dependencies across frames Liu et al. (2021); Lin et al. (2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
The integration of Transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures Liu et al. (2021, 2021). Transformers not only generate semantically accurate captions but also exhibit enhanced coherence and contextual relevance. Their attention mechanism allows for a incorporation of local and global context, enabling a more holistic understanding of the visual content and facilitating caption generation.
The use of Transformers in image and video captioning holds great promise, but there are challenges to be addressed. Efficient fusion of visual features with textual information Radford et al. (2021), handling long-term dependencies in video sequences Liu et al. (2021), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
4.2 Visual Question Answering
Visual Question Answering (VQA) combines computer vision and NLP to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content. Both VQA and VSG involve understanding visual content and generating meaningful responses and the developments in VQA can be directly applied in VSG tasks. Through the use of VQA techniques, VSG models can gain valuable insights into the content and context of images and videos, improving the overall quality of their visual narratives.
Encoder-classifier deep-learning VQA models encode visual and textual inputs and select an answer from a predefined set using a classification process. Image encoders, typically pretrained CNNs, extract visual features from the input images. Question encoders, often RNNs, processes the input questions, converting them into fixed-length latent representations. Fusion techniques combine the visual and textual features, creating a fused representation that captures their interactions that is then classified using fully connected layers, mapping them to possible answer classes. Traditional VQA models are trained using supervised learning with large-scale datasets, where image-question-answer triplets are provided as training examples. The parameters are optimized to minimize the loss between predicted answers and ground truth answers during training. Some examples are BUTD Anderson et al. (2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” Kazemi and Elqursh (2017), that combines LSTMs with multiple attention layers.
The transition from encoder-classifier methods to Transformer methods in VQA was driven by the Transformer’s ability to capture long-range dependencies and global contextual information. Transformers offered the potential to more efficiently capture the interactions between visual and textual modalities, enabling a more comprehensive understanding of the relationship between images and questions. In these VQA models, images and questions are encoded using Transformer encoders and self-attention mechanisms capture the dependencies between words and the visual features, allowing the model to focus on relevant image regions while processing the question. The fused encoded representations are then fed into classification layers to generate answers Bao et al. (2022); Wang et al. (2023, 2022, 2023). Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers Chen et al. (2022).
5 Visual Story Generation
VSG combines computer vision and NLP to create coherent narratives from sequences of images or videos. This task extends beyond simple image captioning by requiring models to understand temporal relationships, infer causal connections between visual scenes, and generate engaging stories that capture the underlying narrative flow. Unlike video captioning, which focuses on objective description of visual content, VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content.
The field has evolved from early encoder-decoder architectures using CNN and RNN to Transformer-based models that leverage large-scale pretraining. These approaches must address challenges including visual understanding, temporal coherence, and narrative creativity while maintaining grounding in the provided visual content. The following subsections present the key methodologies and their contributions to advancing VSG capabilities.
5.1 Visual Storytelling with Convolutional and Recurrent Neural Networks
Huang et al. Huang et al. (2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extended the image captioning technique. It extracts features from each image using a pretrain CNN called AlexNET Krizhevsky et al. (2012). It then encodes image sequences by using a GRU over the features of each individual image. The story decoder model consists of another GRU and generates narratives word by word. Four different decoding methods, shown in Table 3, were considered: beam search, greedy search, greedy search with removal of duplicates, and greedy search with removal of duplicates and incorporation of grounded words.
Method METEOR Score
Beam=10 23.13
Greedy 27.76
Greedy - Duplicates 30.11
Greedy - Duplicates + Grounded 31.42
Table 3: Results of each decoding approach using evaluated using METEOR scores on the VIST Huang et al. (2016) dataset. Red indicates the lowest score, green the highest.
Results indicate that beam search alone does not perform well in generating high-quality stories. This contrasts with the results for image captioning, where beam search is more effective than greedy search. The reason is that generating stories with beam search results in generic high-level descriptions such as “This is a picture of a dog” originating from the label bias problem. On the other hand, by using a beam size of 1 (greedy search) there is an improvement in the quality of the generated stories.
Since stories generated using a greedy search may sometimes contain repeated words or phrases that can adversely affect textual quality, a duplicate-removal heuristic was added to ensure that content words cannot be produced more than once on a given story. Finally, to further improve quality, the previous approach was extended to include “visually grounded” words. These are words that are authorized for use by the story model only if they are licensed by the caption model: for a word to be included in the story, it must have appeared with some frequency in the captions generated by the caption model for the sequence of images. Fig. 1 shows an example of the generated stories for an image sequence.
Figure 1: Sequence of images used to generate stories using different decoding methods. Beam search: This is a picture of a family. This is a picture of a cake. This is a picture of a dog. This is a picture of a beach. This is a picture of a beach. Greedy: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The dog was enjoying the water. The dog was happy to be in the water. Greedy - duplicates: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The kids were playing in the water. The boat was a little too much to drink. Greedy - duplicates + grounded words: The family got together for a cookout. They had a lot of delicious food. The dog was happy to be there. They had a great time on the beach. They even had a swim in the water. Adapted from Huang et al. (2016).
5.2 Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Park et al. Park and Kim (2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet Simonyan and Zisserman (2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence. CRCN also introduces a local coherence model that focuses on maintaining coherence and context within the generated sentence sequences. It also utilizes parse trees to detect grammatical roles and structural attributes, ensuring that the generated sentences integrate with the broader narrative context. During training, correctly aligned image-sentence pairs are given higher scores than misaligned pairs.
The evaluation of CRCN considered several automatic metrics and several baseline works, including one that excludes the entity coherence model. Table 4 shows the results for the NY City Park and Kim (2015) and Disneyland Park and Kim (2015) datasets. Both datasets were introduced by Park et al. Park and Kim (2015) and consist of image sequences with the corresponding blog posts. In these datasets, results were superior to all baselines. The exception, where it was a close second, was with the baseline without entity coherence and in only certain metrics. The work was further validated by user studies conducted via Amazon Mechanical Turk (AMT), showing that human evaluators prefer CRCN-generated sequences over several baselines.
Metric B-1 B-2 B-3 B-4 C M R@1 R@5 R@10
NY City 26.83 5.37 2.57 2.08 30.9 7.69 11.67 31.19 43.57
Disneyland 28.40 6.88 4.11 3.49 52.7 8.78 14.29 31.29 43.20
Table 4: Performance metrics of CRCN on New York City and Disneyland Datasets. Green values represent state-of-the-art results at the time. B stands for BLEU, C for CIDEr, M for METEOR and R for Recall Park and Kim (2015).
Fig. 2 shows an example of a sequence of images and the corresponding generated sentences.
Figure 2: Sequence of images used by CRCN in a story. (1) One of the hallway arches inside of the library (2) As we walked through the library I noticed an exhibit called lunch hour nyc it captured my attention as I had also taken a tour of nyc food carts during my trip (3) Here is the top of the Chrysler building everyone’s favorite skyscraper in new york. (4) After leaving the nypl we walked along 42nd st. (5) We walked down fifth avenue from rockefeller centre checking out the windows in saks the designer stores and eventually making our way to the impressive new york public library. Adapted from Park and Kim (2015).
In addition to the work by Park et al. Park and Kim (2015), which was the first to use a bidirectional RNN for visual storytelling, other works have also employed bidirectional RNNs architectures. XE Wang et al. (2018) and AREL Wang et al. (2018) have proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
5.3 Transitional Adaptation of Pretrained Models for Visual Storytelling
Yu et al. Yu et al. (2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks. It aims at bridging the gap between pretrained language models and visual encoders. In contrast with previous models for vision-to-language generation tasks, which typically pretrain a visual encoder and a language generator separately and then jointly fine-tune them for the target task, TAPM proposes a transitional adaptation task to harmonize the visual encoder and language model for downstream tasks like VSG, mitigating discord between visual specificity and language fluency arising from separate training on large corpora of visual and text data. TAPM introduces an approach that adapts multimodal modules through a simpler alignment task focusing solely on visual inputs, eliminating the need for text labels.
TAPM components aim at enhancing the quality of textual descriptions for visual content in storytelling tasks. They are a visual encoder, a language generator, adaptation training, sequential coherence loss, training with adaptation loss, and a fine-tuning and inference process. The visual encoder, a pretrained model, extracts features from images or videos. In TAPM, it becomes integral during the adaptation phase, where it integrates with the language generator to fuse visual and textual information. The pretrained language generator model is responsible for converting visual information into textual descriptions. During the adaptation phase, it generates video and text embeddings, aligning textual representations with the corresponding visual features, based on a sequential coherence loss function.
The loss function divides sequential coherence into three components: past, current, and future matching losses. The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video i1i-1 and distancing it from those of non-sequential videos. The future matching loss projects $b_{s_{i}}$ via a distinct FC layer $f_{f}$, aligns with the subsequent visual representation $v_{b_{i+1}}$. The current matching loss then aligns the current visual representation $v_{b_{i}}$ with $b_{s_{i}}$ through another FC layer $f_{c}$. These components are unified by the FC layer projections in their respective visual spaces, pulling the embeddings of correct matches closer and pushing incorrect matches further apart. Margin ranking losses are utilized to implement this concept, contrasting correct matches with incorrect ones. The final sequential coherence loss for a given video $i$ is formulated as shown in Eq. 1, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
$L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$ (1)
$+$ $\sum_{j\neq i}\max(0,1+\cos(v_{b_{j}},f_{c}(b_{s_{i}}))-\cos(v_{b_{i}},f_{c}(b_{s_{i}})))$
$+$ $\sum_{j\neq i+1}\max(0,1+\cos(v_{b_{j}},f_{f}(b_{s_{i}}))-\cos(v_{b_{i+1}},f_{f}(b_{s_{i}}))),$
TAPM uses a split-training strategy to optimize model performance. Initially, the visual encoder undergoes adaptation training with the adaptation loss, while the text encoder and language generator remain fixed. Subsequently, all components are jointly updated with the generation loss, allowing the model to optimize the adaptation task before addressing the more challenging generation objective. After the adaptation and split-training phases, TAPM undergoes fine-tuning. The model is then ready for the inference phase, generating captions for unseen visual inputs.
Table. 6 and 6 shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets. Table. 7 show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019. On VIST it surpasses the XE Wang et al. (2018) and AREL Wang et al. (2018) baselines in relevance, expressiveness, and concreteness as shown in Table 8. These results highlight the strengths in word choice and contextual accuracy, showcasing its ability to capture causal relationships between images. However, the score is still far from human performance, indicating that there is still room for improvement.
Models C M R
Huang et al. Huang et al. (2016) - 31.4 -
AREL Wang et al. (2018) 9.4 35.0 29.5
StoryAnchor Modi and Parde (2019) 9.9 35.5 30.0
HSRL Huang et al. (2018) 10.7 35.2 30.8
INet Jung et al. (2020) 10.0 35.6 29.7
TAPM 13.8 37.2 33.1
Table 5: Quantitative results on the VIST test set. C stands for CIDEr, M for METEOR, and R for ROUGE-L. Green indicates the highest score, red the lowest.
Models Public Test Blind Test
C R C R
Baseline Park et al. (2018) 7.0 12.0 6.9 11.9
XE Wang et al. (2018) 7.2 11.5 - -
AREL Wang et al. (2018) 7.3 11.4 - -
TAPM 10.0 12.3 8.8 12.4
Table 6: Quantitative results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021).
Models Scores
Human 1.085
Official Baseline Park et al. (2018) 4.015
TAPM 3.670
Table 7: Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021).
TAPM vs XE TAPM vs AREL
Choice (%) TAPM XE Tie TAPM AREL Tie
Relevance 59.9 34.1 6.0 61.3 32.8 5.9
Expressiveness 57.3 32.3 10.4 57.3 34.0 8.7
Concreteness 59.1 30.3 10.7 59.6 30.4 10.0
Table 8: Human evaluation results on VIST. Higher is better. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021)
5.4 Interactive and Creative Visual Storytelling
Building upon the foundational CNN and RNN approaches, interactive and creative visual storytelling represents an important evolution in the field, introducing intermediate approaches that bridge visual comprehension with narrative creativity. These methods combine elements of image captioning and VQA while introducing narrative structure and creative interpretation, addressing limitations of purely technical approaches.
Lukin et al. Lukin et al. (2018) introduced a three-module pipeline for creative visual storytelling that systematically approaches the challenge of generating narratives from visual content. Their pipeline consists of object identification, single-image inferencing, and multi-image narration, that serve as a preliminary design for building a creative visual storyteller. The approach defines computational creative visual storytelling as one with the ability to alter the telling of a story along three aspects: to speak about different environments, to produce variations based on narrative goals, and to adapt the narrative to the audience. This modular approach demonstrates how the foundational capabilities of image captioning and VQA can be systematically extended and combined to support more complex narrative generation tasks. The pipeline explicitly separates visual understanding from narrative construction, allowing for more systematic development and evaluation of each component while maintaining the overall coherence required for storytelling.
Interactive visual storytelling has also explored user engagement and personalization in narrative generation. Halperin and Lukin Halperin and Lukin (2023) examined how creative visual storytelling can serve as an anthology for narrative intelligence, investigating the intersection of human creativity and automated story generation through an analysis of 100 visual stories from authors who participated in a systematic creative process of improvised story building based on image sequences. Their work on surreal visual storytelling Halperin and Lukin (2024) further explores how visual narrative systems can handle ambiguous or dreamlike imagery, investigating AI “hallucination” by stress-testing a visual storytelling algorithm with different visual and textual inputs designed to probe dream logic inspired by cinematic surrealism. These approaches highlight the importance of modularity in visual storytelling systems, where different components can be optimized independently while contributing to the overall narrative coherence. The integration of creative elements with systematic visual analysis provides a foundation for more sophisticated VSG systems that can balance factual visual description with imaginative narrative construction.
5.5 Knowledge-Enhanced Visual Storytelling
A challenge in VSG is the tendency of end-to-end approaches to produce monotonous stories with repetitive text and limited lexical diversity. This limitation arises because existing approaches are constrained by the vocabulary and knowledge available in single training datasets. To address this challenge, researchers have explored knowledge enhanced approaches that leverage external resources to enrich the story generation process.
Hsu et al. Hsu et al. (2020) introduced KG-Story, a three-stage framework that allows VSG systems to take advantage of external knowledge graphs to produce more diverse stories. The framework implements a distill-enrich-generate approach: first distilling a set of representative words from input prompts, then enriching the word set using external knowledge graphs, and finally generating stories based on the enriched word set. This framework allows the use of external resources not only for the enrichment phase, but also for the distillation and generation phases. The KG-Story framework operates through three distinct stages. In stage 1, an image-to-term model distills representative terms from each input image, creating conceptual representations that capture the essential elements of the visual content. In stage 2, external knowledge graphs are utilized to identify possible connections between the extracted term sets from different images, generating enriched term paths that capture relationships and associations not explicitly visible in the images. In stage 3, a Transformer architecture transforms these term paths into coherent stories, incorporating techniques such as length difference positional encoding and repetition penalties to improve narrative quality. Evaluation results demonstrate that stories generated by KG-Story are on average ranked better than previous state-of-the-art systems in human ranking evaluations. The approach successfully addresses the vocabulary limitations of traditional end-to-end methods while maintaining narrative coherence.
Interactive approaches to knowledge enhanced storytelling have also been explored. Hsu et al. Hsu et al. (2019) introduced Dixit, an interactive visual storytelling system that allows users to iteratively compose stories through term manipulation. The system extracts text terms describing objects and actions from photos, then allows users to add new terms or remove existing ones before generating stories based on these modified term sets. Behind the scenes, Dixit uses an LSTM model trained on image caption data to distill terms from images and utilizes a Transformer decoder to compose context coherent stories. This approach opens up possibilities for interpretable and controllable visual storytelling, allowing users to understand the story formation rationale and to intervene in the generation process.
These knowledge enhanced approaches represent an important advancement in VSG, demonstrating how external knowledge resources can be systematically integrated to overcome the limitations of purely data-driven methods. By explicitly modeling the relationship between visual content and broader conceptual knowledge, these systems can generate more diverse, engaging, and contextually rich narratives while maintaining controllability and interpretability.
5.6 Iterative and Planning-Based Approaches
Traditional VSG models typically generate stories in a single forward pass. However, creative writers use their knowledge and worldview to put disjointed elements together to form a coherent storyline, and work and rework iteratively toward perfection. VSG models, however, make poor use of external knowledge and iterative generation when attempting to create stories. This observation has motivated the development of iterative and planning-based approaches that more closely mirror human creative writing processes. Hsu et al. Hsu et al. (2021) introduced PR-VIST (Plot and Rework Visual Storytelling), a framework that represents the input image sequence as a story graph in which it finds the best path to form a storyline. PR-VIST then takes this path and learns to generate the final story via a re-evaluating training process. The framework draws inspiration from the human creative writing process, which involves plotting (planning the overall narrative structure) and reworking (iteratively refining the story content).
The PR-VIST framework operates through two main stages that correspond to the plotting and reworking phases of creative writing. In stage 1 (story plotting), the system constructs a story graph from the input image sequence using external knowledge graphs, including VIST and Visual Genome knowledge graphs. A storyline predictor model identifies the best path through this graph to form a coherent storyline, creating a structured narrative plan that connects the images through meaningful conceptual relationships. This plotting stage essentially creates a roadmap for the story that will be generated. In stage 2 (story reworking), the framework generates the actual story text based on the predicted storyline path. The approach uses a length controlled Transformer that is first pre-trained on the ROC Story dataset and then fine-tuned on the VIST dataset with a discriminator component to encourage higher quality story generation. The re-evaluating training process allows the model to iteratively improve the story generation by learning from feedback about narrative quality and coherence.
Evaluation results demonstrate that this framework produces stories that are superior in terms of diversity, coherence, and humanness, per both automatic and human evaluations. An ablation study shows that both plotting and reworking contribute to the model’s superiority. The explicit separation of planning and generation phases allows for better control over narrative structure while maintaining the flexibility needed for creative story generation. This planning based approach represents an important advancement toward more human-like story generation, demonstrating how computational systems can benefit from explicitly modeling the structured creative processes that human writers naturally employ.
5.7 Combining LLMs (GPT-2) and Character Features
Hong et al. Hong et al. (2023) proposed CharGrid (Character-Grid Transformer), a Transformer model that integrates diverse input features, from global image attributes to character-specific details. It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s Liu et al. (2021), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing. The trainable image and character encoders process the image feature inputs. The character grid is flattened and fed to a FC layer. The Transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained Generative Pre-trained Transformer (GPT)-2 tokenizer. The model architecture is shown in Fig. 3.
Figure 3: CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from Hong et al. (2023).
During training, model parameters, excluding the vision backbone, are randomly initialized. The training objective involves maximizing the likelihood of image sequence-story pairs through backpropagation. Nucleus sampling Holtzman et al. (2019) is employed for validation, and the METEOR score is used for evaluation.
CharGrid has two additional variants: ObjGrid (Object-Grid Transformer) and EntiGrid (Entity-Grid Transformer). ObjGrid replaces character features with object features. This model incorporates global features and object features, with an object grid representing coherence based on the similarity between global image features and object features. It is similar to CharGrid but includes both character and object features. EntiGrid incorporates global, character, and object features, with an entity grid representing coherence through the similarity between global image features and both character and object features. These variants explore the impact of different input feature combinations on model performance. CharGrid, with its emphasis on character coherence, serves as the primary model, while ObjGrid and EntiGrid address the contributions of object-related features.
The results in Table 9 show that CharGrid outperforms TAPM with character features and GPT-2 with character features on various metrics, emphasizing the effectiveness of character grid representations for coherence in VSG. A crowd-sourcing experiment was conducted with 28 workers to obtain binary judgments on grammaticality, coherence, diversity, and visual groundedness of the generated stories. The results in Table 10 show that TAPM with character features excels in visual groundedness over plain TAPM, while CharGrid surpasses TAPM with character features across all metrics. Two-sided binomial tests support that the character grid representation yields superior stories, affirming the reference-based metrics findings.
Model B-1 B-2 B-3 B-4 M R-L C
GPT-2 38.65 20.28 9.78 4.68 31.64 24.24 1.66
GPT-2 + obj 40.65 21.35 10.2 4.87 31.69 24.05 1.85
GPT-2 + char 39.95 21.04 10.11 4.92 31.85 24.19 1.57
GPT-2 + obj,char 40.41 21.44 10.56 5.06 32.03 24.38 1.87
TAPM 39.85 21.7 10.72 5.19 32.38 25.09 1.48
TAPM + obj 40.86 22.13 10.83 5.25 32.34 24.91 1.82
TAPM + char 40.03 21.68 10.66 5.18 32.42 24.88 1.4
TAPM + obj,char 40.87 21.99 10.72 5.06 32.48 24.87 1.59
ObjGrid + obj 47.66 25.26 11.95 5.42 32.83 24.42 4.68
EntityGrid + obj,char 45.83 24.85 12.11 5.7 32.68 24.89 3.53
CharGrid + char 47.71 25.33 11.95 5.42 33.03 25.01 4.83
Table 9: Results using different input features on the test set of VWP. Global features are always included. “+ obj” and “+ char” represent that object and character features are also included, respectively. BLEU (B), METEOR (M), ROUGE-L (R-L), and CIDEr (C) values are averages of three runs with different random seeds. Green indicates the highest score, red the lowest. Adapted from Hong et al. (2023).
Model G C VG D
TAPM + char vs. TAPM +2.45 +1.99 +3.99* +1.69
CharGrid vs. TAPM + char +6.49** +8.41** +6.25* +11.06**
Table 10: Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality (G), Coherence (C), Visual Groundedness (VG), and Diversity (D). The numbers are percentages. * means p-value << 0.05. ** means p-value << 0.01. Adapted from Hong et al. (2023).
5.8 Visual Story Generation using Graphs for Event Ratiocination
Zheng et al. Zheng et al. (2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters. The work addresses several downstream tasks, including VSG. The work introduces a framework to improve the integration and interpretation of visual and textual data, using hypergraphs to address challenges in multimodal interactions and temporal dynamics. Fig. 4 shows this architecture.
Figure 4: Architecture of HEGR by Zheng et al. Adapted from Zheng et al. (2021).
The method begins with feature extraction: visual features are extracted using Mask R-CNN He et al. (2017) to detect objects and their spatial characteristics within images. Textual features are extracted from associated descriptions or captions using BERT Devlin et al. (2019), capturing semantic information. In the second phase, semantic graphs are constructed for each modality: nodes represent detected features or semantic tokens, while edges depict relationships within the same modality. To capture dependencies between modalities and across various temporal segments, hypergraphs are employed, linking semantic graphs based on contextual and temporal relevance, forming a high-dimensional interaction map. In the third phase, graph convolution is used to update node representations within each graph, allowing contextual information aggregation through learned adjacency matrices. Hyper-graph convolution extends this to include higher-order relationships across multiple graphs and temporal instances. In the fourth phase, a self-attention mechanism refines graph representations by dynamically adjusting the significance of nodes within and across modalities. This process focuses on the most relevant features for narrative generation. Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART Lewis et al. (2019).
The method was evaluated on the VIST dataset, achieving state-of-the-art performance in terms of BLEU-4, ROUGE-L and METEOR scores achieving 16.4, 37.4, and 37.8, respectively.
5.9 Discussion
The works presented in this section have made substantial contributions to the field of VSG, establishing robust benchmarks for the assessment of emerging models and methodologies.
Specifically, the first three techniques, presented in Sections 5.15.2, and 5.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation. The methods detailed in Sections 5.1 and 5.3, offered a comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively. Notably, the latter technique surpasses the former, which can be attributed to its transitional training approach.
The study outlined in Section 5.7, CharGrid, does not present findings on the VIST dataset. However, this method is benchmarked against the method detailed in Section 5.3, TAPM, on the VWP dataset, where it demonstrates superior performance across all evaluated metrics. This comparison is shown in Table 9, specifically within the rows labeled “TAPM + obj,char” and “CharGrid”. This method differentiates itself from the previous by focusing on the extraction of object-level features from visual inputs rather than employing a CNN to encode the entire visual input. Furthermore, it leverages the capabilities of an LLM, using its extensive world knowledge to enhance language modeling, illustrating a novel integration of visual and linguistic elements in story generation.
Finally, the method discussed in Section 5.8, HEGR, introduces a different approach that uses hypergraphs to model the contextual and temporal relationships between visual and textual data. This approach is explicitly able to model some story elements, namely the characters and the setting.
Table 11 summarizes the performance of the methods described above, along with other methods from the literature that contributed to the field of VSG. The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN Ren et al. (2015) and Mask-RCNN He et al. (2017), and LLMs like GPT-2 Radford et al. (2019) and BART Lewis et al. (2019). The first two methods by Park et al. (2015) Park and Kim (2015) and Park et al. (2019) Park et al. (2018) are not comparable to the other methods as they do not use the VIST dataset. The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) Hong et al. (2023) that is evaluated on the VWP dataset. The authors of CharGrid trained and evaluated the performance of TAPM on the VWP Hong et al. (2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset. It is however not possible to know if CharGrid would perform better than HEGR on the VIST dataset without further experiments.
Model Encoding Decoding Dataset B-4 R-L M
Park et al. (2015) Park and Kim (2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
Park et al. (2019) Park et al. (2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
Huang et al. (2015) Huang et al. (2016) CNN, GRU GRU VIST - - 31.4
Xu et al. (2015) Xu et al. (2015) VGGNet LSTM VIST - 28.94 32.98
Yu et al. (2017) Yu et al. (2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
CST (2018) Gonzalez-Rico and Pineda (2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
AREL (2018) Wang et al. (2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
HSRL (2019) Huang et al. (2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
GLACNet (2019) Kim et al. (2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
StoryAnchor (2020) Modi and Parde (2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
INET (2020) Jung et al. (2020) CNN, GRU GRU VIST 14.7 29.7 35.5
AOG-LSTM (2023) Liu et al. (2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
CoVS (2023) Gu et al. (2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
TAPM (2021) Yu et al. (2021) 1 Faster R-CNN, ResNet-152 GPT-2 s VIST - 33.1 37.2
HEGR (2021) Zheng et al. (2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
TAPM (2021) Hong et al. (2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
CharGrid (2023) Hong et al. (2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
Table 11: Analysis of Visual Story Generation works. The columns “B-4”, “R-L”, and “M” represent the BLEU-4, ROUGE-L, and METEOR scores, respectively. All works, except Huang et al. (2015) Huang et al. (2016) and CST (2018) Gonzalez-Rico and Pineda (2018), use attention mechanisms. All works use images as input, except ActivityNet Fabian Caba Heilbron and Niebles (2015) (it uses videos). The mark 1 indicates works that used extra training data (direct comparison may not be fair). The mark 2 indicates works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
6 Real-World Applications
As VSG technologies mature, they are increasingly being deployed in real-world scenarios. These applications demonstrate the practical potential of VSG systems and highlight new challenges that arise when deploying these technologies in dynamic, context-rich environments. This section explores emerging applications that integrate visual storytelling with contextual information, location awareness, and real-time data processing.
6.1 Location-Aware Visual Storytelling
Suwono et al. Suwono et al. (2023) introduced location-aware visual question generation with lightweight models, representing an important advancement in spatially-grounded narrative applications. Their work addresses the challenge of generating contextually relevant questions and narratives that are informed by geographical location, opening new possibilities for travel documentation, educational applications, and location-based social media platforms. The location-aware approach integrates geographical metadata with visual content analysis to create narratives that are visually grounded and spatially contextualized. This integration enables the generation of stories that can reference specific landmarks, cultural contexts, and geographical features that would be meaningful to users familiar with particular locations. The use of lightweight models addresses practical deployment constraints, making the technology suitable for mobile applications and resource-constrained environments.
6.2 Automotive and Real-Time Context Integration
Belz et al. Belz et al. (2024) explored story-driven approaches in automotive contexts, investigating how real-time contextual information can enhance automated storytelling systems. Their research represents a step toward practical deployment of VSG in everyday scenarios, particularly in automotive systems where visual storytelling could enhance the driving experience and provide meaningful documentation of journeys. The automotive application domain presents challenges and opportunities for visual storytelling. In-car camera systems continuously capture visual data, creating an opportunity for real-time narrative generation that could help drivers and passengers understand and remember their journeys. The integration of real-time context information—such as speed, location, weather conditions, and traffic patterns enables the generation of richer, more informative narratives that extend beyond visual description. Their work demonstrates how dynamic contextual data can be incorporated into storytelling algorithms, showing measurable improvements in narrative relevance and user engagement when real-time information is provided. This approach could lead to applications in autonomous vehicles, where passengers might receive automatically generated travel narratives, or in fleet management systems where journey documentation becomes an automated process.
7 Conclusions
This survey reviewed the field of VSG, explaining its development and current methods. It started by explaining basic story elements, datasets and evaluation metrics used to judge the quality of the stories. The discussion pointed out the need for better metrics to assess complex tasks like VSG accurately. The document also covered related areas like image and video captioning and VQA, discussing how improvements in these fields have helped enhance VSG. The discussion then moved on to the evolution of VSG methods, starting with early approaches that used CNNs and RNNs to methods based on LLMs and hypergraphs. Additionally, we examined emerging real-world applications that demonstrate the practical potential of these technologies in domains such as automotive systems and location-aware storytelling.
Despite the advances, challenges remain for VSG. In the following subsections, we discuss some of the key challenges and opportunities in the field that could guide future research directions. We start with the ones that are more concrete and could influence the field in the short term and end with the ones that are more abstract and could potentially influence the field in the long term.
7.1 Exploring Large Language Models
The field of VSG using LLMs is under-explored, with only a few works delving into this domain. To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Qwen3-235B-A22B Yang et al. (2025) which is the best performing open-source model on Chatbot Arena or Gemini 2.5 Pro Comanici et al. (2025) which is the best performing proprietary model. The scarcity of the integration of LLMs with visual elements presents an opportunity for further research, exploring the potential of these models in generating more coherent and contextually relevant stories.
7.2 Automatic Evaluation Metrics
The need for robust and efficient automatic evaluation metrics is emphasized in our survey. While existing metrics like METEOR and CIDEr have been widely adopted, they do not fully capture the nuances of story quality. The metrics mostly focus on the similarity between the generated text and the reference text. A story can be retold in various ways, employing different perspectives while preserving the original message. Consequently, a generated text might deviate significantly from the reference text in terms of specific wording and still deliver the same story in a coherent way. Thus, while these metrics provide valuable insights into certain dimensions of text generation, they did not fully encompass the multifaceted nature of storytelling. As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al. (2023) and so metrics to evaluate stories could be developed based on LLMs. We advance that these metrics could be based on the existing systematic scoring systems Zheng et al. (2023) that were developed to evaluate the performance of LLM-based chat bots.
7.3 Decomposition of Story Generation
Addressing the complexity of stories, we underscore the importance of decomposing the story generation process into modular components. Each module can focus on specific attributes such as author goals, character interactions, consistency, etc. This decomposition has already been explored with modeling characters in CharGrid Yu et al. (2021) and the modeling of characters and setting in HEGR Zheng et al. (2021). We believe that explicit modeling of other story elements and abstract concepts such as emotions and culture, possibly using hypergraphs or similar structures could further enhance the quality of generated stories.
7.4 Hybrid Systems and Knowledge Integration
Building on the decomposition concept, hybrid systems, combining planning-based techniques and machine learning approaches, present an intriguing direction. Hyper-graphs, as seen in HEGR Zheng et al. (2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems. The integration of multiple knowledge sources can enrich the diversity and quality of generated stories, paving the way for more robust and context-aware narrative creation.
7.5 Real-World Deployment and Context Integration
The transition from academic research to practical applications represents a crucial next step for VSG technologies. Recent work in location aware storytelling and automotive applications demonstrates the potential for deploying these systems in real-world scenarios. Future research should focus on developing robust systems that can integrate multiple contextual information sources including location, temporal data, and real-time sensor information while maintaining computational efficiency for deployment in resource constrained environments such as mobile devices and embedded systems. The success of these real-world applications will ultimately determine the broader impact of VSG research on society.
Acknowledgements.
[…]